# Federated Automated Feature Engineering

### v1

@@ 4.1 Horizontal Setting
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, denote the list of feature strings of engineered features on client $k$ as $\mathcal{F}_{k}$, AutoFE routine $A_{k}:\{X_{k},y_{k}\}\rightarrow\mathcal{F}_{k}$ for each client $k$
@@
3 for each j=1,…,J do
4       for each client $k=1,2,...,Q$ do
5             $\mathcal{F}_{k}$ = $A_{k}(\{X_{k},y_{k}\})$
6             Send $\mathcal{F}_{k}$ to server
7      $\mathcal{F}$ = $\mathcal{F}_{1}\cup\mathcal{F}_{2}\cup...\cup\mathcal{F}_{Q}$
@@ 11             $\mathbb{L}=\{\;\}$
18      Find best score and corresponding selection mask $g_{best}$ across all sparsity-refinement iterations
@@ 19       $\mathcal{F}=g_{best}\odot\mathcal{F}$
20       Each client builds new engineered numerical features $X_{k}$ from $\mathcal{F}$ and previous $X_{k}$.
@@
Algorithm 1 Horizontal-FLAFE
@@ 3       Draw random sparsity $\mathcal{S}_{c}$ from Uniform($s_{a}$,$s_{b}$)
We first describe our algorithm for the horizontal FL setting; as far as we know, this is the first horizontal federated AutoFE algorithm to be developed. First, AutoFE is performed on each client separately only with local data on that client. This allows great flexibility in selecting an AutoFE algorithm, as the structure of Horizontal-FLAFE allows for any choice. Then, each client sends the feature string representation of the engineered features to the server. It is important to note that no data is actually sent to the server, only the string representation of the engineered feature such as log(X001**2/X002). The server then finds the union of all of the sets of engineered feature strings from each client. The union of the engineered feature strings is sent back to each client, where the clients then parse the feature string and compute the actual numerical feature vector.
@@
The next step of the horizontal algorithm is selecting the best engineered features in a federated fashion. Our approach is inspired by Hyperband [hyperband] where successive halving is used with varying resource allocations (in this case the resource is communications between client and server). First, a large set of random candidate feature masks are generated. The feature masks are vectors of 0s and 1s that determine which features to keep and which to remove. The operation of masking the feature vector is denoted as $g\odot\mathcal{F}$. For each of these feature masks, each client removes the features specified by the feature mask and then the model is trained and evaluated using FedAvg (or any other horizontal FL algorithm) but with only a very limited number of communication rounds. Then, the top performing feature masks are kept, and the rest discarded. The next round repeats the same process but with a larger number of communication rounds in the FL training. This is repeated until the best performing masks are found with a large number of communication rounds.
@@
This process is repeated for various number of levels similarly to Hyperband. Finally, we cannot simply uniformly draw our selection masks because this results in an average sparsity of 0.5. However, we find that depending on the dataset, base AutoFE algorithm used, and degree of non-IID across clients, the best performing sparsity on the feature mask greatly differs from 0.5. Thus, we have an outer loop that gradually refines the best performing sparsity. Finally, this entire process is repeated in an outermost loop in order to generate increasingly complex features as combinations of previously engineered features. However, the outputs from AutoFE are generally already fairly complex, so the number of outermost iterations is usually kept to a low number.
Figure 1: Workflow of one outer iteration of Horizontal-FLAFE with 3 clients and just a few engineered features. Typically there are more clients and each local AutoFE process generates many more complicated engineered features.
4.2 Vertical Setting
The basic idea for the vertical algorithm is that for features that are on separate clients and can’t be directly shared, we need to homomorphically encrypt the vectors, combine the features in the encryped space, encrypt and then add Laplace noise, and finally send the combined features back to the clients to decrpyt. This combined feature can then be tested to see if it is a useful new feature.
The only homormorphic encryption schemes that are fast enough for practical use preserve addition and multiplication operations. Thus, any complicated functions must be decomposed into a Taylor series of additions and multiplications. However, since homomorphic encryption is very costly, especially for large datasets, we want to minimize the number of times we need this operation. This is why we first find the feature importance of each feature held on the client, and only the most important features on this client are combined with the most important features on other clients.
Once all of the engineered features are created and sent back to clients, the best engineered features are selected in the same way that the feature selection mask is found in Algorithm 1. The successive halving steps require training and evaluating in the vertical FL setting, so vertical FL training algorithms are needed. It is important to note that most homomorphic encryption schemes require specifying the downstream function while encrypting the inputs, while in our approach we apply various functions (combinations of multiplications and additions) to the encrypted inputs. Fortunately, BFV homomorphic encryption does not require the specification of the downstream function while encrypting and fits well within our approach [bfv1][bfv2]. This process is shown in Figure 2.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N}$ belonging to client $k$, BFV Homomorphic encryption function $H(\cdot)$
2 $\mathcal{F}_{final}=[\;]$
3 for each outer iteration $i=1,...,O$ do
4       for each client $k=1,...,Q$ do
5             $\mathcal{F}_{k}=[\;]$
[…]
7             $\mathcal{H}_{k}=[\;]$
8             for each $f\in\mathcal{F}_{important,k}$ do
9                   Homomorphic encrypt feature vector denote as $H(f)$
10                   $\mathcal{H}_{k}=\mathcal{H}_{k}\cup H(f)$
12      Server executes CreateAndSendNonLocalCombinations()
13       Clients receive new combined features from server
14       Each client decrypts the features they were sent $F_{i,j}=\text{Dec}(H_{i,j})$ back to the original feature space
15       Select best newly engineered features, denoted as $\mathcal{F}_{new}$, using Lines 6-14 of Algorithm 1, this requires Vertical FL training algorithms
16       $\mathcal{F}_{final}=\mathcal{F}_{final}\cup\mathcal{F}_{new}$
Return $\mathcal{F}_{final}$
Algorithm 3 Vertical-FLAFE
1 Input: features F where $i$-th column is indexed as $F_{i}$, set of candidate bivariate transformations
2 $F_{out}=[\;]$
3 for each $(F_{i},F_{j})\in F\times F$ do
4       for each candidate bivariateFunction() do
5             $F_{out}=F_{out}\cup\text{bivariateFunction}(F_{i},F_{j})$
6
Return $F_{out}$
Algorithm 4 CreateLocalCombinations()
1 Input: Lists of encrypted feature vectors for each client denoted as $\mathcal{H}_{k}$, set of candidate bivariate transformations, $b$ is the scale parameter used for the Laplace noise
2 $\mathcal{H}=\mathcal{H}_{1}\cup...\cup\mathcal{H}_{Q}$
3 $H_{out}=[\;]$
4 for each $(H_{i},H_{j})\in\mathcal{H}\times\mathcal{H}$ do
5       for each candidate bivariateFunction() do
             $H_{i,j}=\text{bivariateFunction}(H_{i},H_{j})+H(\text{Lap}(0|b)^{N})$ // inject Laplace noise
6             Select client that owns feature $i$ or feature $j$ at random with equal probability of selecting between the two
7             Send $H_{i,j}$ back to selected client
8
Algorithm 5 CreateAndSendNonLocalCombinations()
4.3 Hybrid Setting
We make some assumptions on the way that data is gathered across clients in the hybrid setting. While work such as [hyfdca] allows fairly flexible partitions of data as long as each client hold rectangular patches of the overall data, we enforce more strict rules on the way data is split. Specifically, we require that every sample be partitioned across clients in the same way. Essentially, the vertical divisions must run across the entire sample space without any jagged partitions. An example of a suitable partition is shown in Figure 3. Algorithm 6 follows very similarly to Algorithm 6, but has some key differences that are needed in the hybrid setting. First, the process of finding the most important features on each client to combine with other client’s features needs to be done across several clients because each feature is held across many clients in the hybrid setting. This likely means training with horizontal FL across clients that share the same features but different samples to determine feature importances. Furthermore, selection of the best generated features requires hybrid FL training algorithms, of which there are very few and these typically have assumptions on the type of models they work for. For example, as far as we know, there are no hybrid FL algorithms for tree-based models.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, BFV Homomorphic encryption function $H(\cdot)$
2 $\mathcal{F}_{final}=[\;]$
3 for each outer iteration $i=1,...,O$ do
4       for each client $k=1,...,Q$ do
5             $\mathcal{F}_{k}=[\;]$
[…]
6             Find $K$ most important features on client $k$ denoted as $\mathcal{F}_{important,k}\subseteq\mathcal{F}_{k}$, perform this across all clients that contain the specified feature (may require FL training depending on chosen feature selection method)
7             $\mathcal{H}_{k}=[\;]$
8             for each $f\in\mathcal{F}_{important,k}$ do
9                   Homomorphic encrypt feature vector denote as $H(f)$
10                   $\mathcal{H}_{k}=\mathcal{H}_{k}\cup H(f)$
12      Server executes CreateAndSendNonLocalCombinations()
13       Clients receive new combined features from server
14       Each client decrypts the features they were sent $F_{i,j}=\text{Dec}(H_{i,j})$ back to the original feature space
15       Select best newly engineered features, denoted as $\mathcal{F}_{new}$, using Lines 6-14 of Algorithm 1, this requires Hybrid FL training algorithms
16       $\mathcal{F}_{final}=\mathcal{F}_{final}\cup\mathcal{F}_{new}$
Return $\mathcal{F}_{final}$
Algorithm 6 Hybrid-FLAFE
5 Experiments
We conduct basic experiments to evaluate how well Horizontal-FLAFE performs compared to standard AutoFE in the centralized case. We perform experiments on the OpenML586 and Airfoil datasets, the base AutoFE algorithm used is AutoFeat with 2 feature engineering steps. The downstream model used is Lasso with a regularization constant of 0.01. We set $J=1$, $I=1$, $H=1$, and $\mathcal{C}_{1,1}=16$. We split the data across 8 clients in an IID fashion. The clients hold 60% of the overall dataset, and the server holds 20% as a final hold-out test set, and 20% as a validation set used in determining the best feature mask. The models are trained in a federated fashion using FedAvg.
We compute the scores for AutoFE with fully centralized data and observe how close the FL approaches get to this score. Finally, we also compute the test scores for the case with no feature selection mask (the full union of engineered features is used), in order to justify the process of selecting the best feature selection mask.
Figure 4 shows the results on these two datasets. The plot shows the baseline test score, where the model is trained centrally and only the original, raw features are used. The centralized test score is such that AutoFE is performed centrally, and then, the final model is trained centrally with the best engineered features. The “without mask” bar references the final model trained with FedAvg, where all of the engineered features found on each client are used without performing any selection. Finally, the Horizontal-FLAFE bar shows the test score with the final model trained with FedAvg and the best engineered features created and selected using Algorithm 1. It is important to note that we do not compare with any other Horizontal federated AutoFE approaches, because as far as we know, no other horizontal federated AutoFE algorithms exist.
These results demonstrate that Horizontal-FLAFE approaches the same test scores as if AutoFE is performed centrally, and then, the final model is trained centrally. In fact, the test score for Airfoil is higher for Horizontal-FLAFE than centralized AutoFE. We believe this is due to the fact that Horizontal-FLAFE creates many more engineered features as the union is taken across all clients, and we do not expect this to hold for non-IID experiments. Furthermore, there is a clear increase in test score between the optimal selection mask and without a selection mask (using the full union of engineered features); this justifies our resource-aware selection mask optimization. The best feature mask for OpenML586 had a sparsity of 0.902 and for Airfoil the best sparsity was 0.708, indicating that our sparsity refinement approach is better than just uniformly randomly selecting the candidate feature masks.
Figure 4: Performance of Horizontal-FLAFE on two different datasets.
6 Conclusions and Future Work
We contribute three AutoFE algorithms that operate in the three key federated learning settings: horizontal, vertical, and hybrid. We provide experimental evidence that Horizontal-FLAFE performs similarly to performing the AutoFE process in a centralized manner. Our future work is focused on expanding the experimental results across more datasets and choices of algorithm hyperparameters and having extensive experiments performed in the vertical and hybrid FL settings.
[…]

### v2

@@ 7.3 Alternative Algorithms
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, denote the list of features as $\mathcal{F}$,
@@ 6       $\mathbb{L}=\{\;\}$
13Find best score and corresponding selection mask $g_{best}$ across all sparsity-refinement iterations
@@ 14 $\mathcal{F}=g_{best}\odot\mathcal{F}$
Algorithm 5 Fed-Halving: Federated Feature Selection
@@ 3       Draw random sparsity $\mathcal{S}_{c}$ from Uniform($s_{a}$,$s_{b}$)
We propose a federated feawture selection algorithm in Algorithm 5. Our approach is inspired by Hyperband [hyperband] where successive halving is used with varying resource allocations (in this case the resource is communications between client and server). First, a large set of random candidate feature masks are generated. The feature masks are vectors of 0s and 1s that determine which features to keep and which to remove. The operation of masking the feature vector is denoted as $g\odot\mathcal{F}$. For each of these feature masks, each client removes the features specified by the feature mask and then the model is trained and evaluated using FedAvg (or any other horizontal FL algorithm) but with only a very limited number of communication rounds. Then, the top performing feature masks are kept, and the rest discarded. The next round repeats the same process but with a larger number of communication rounds in the FL training. This is repeated until the best performing masks are found with a large number of communication rounds.
@@
This process is repeated for various number of levels similarly to Hyperband. Finally, we cannot simply uniformly draw our selection masks because this results in an average sparsity of 0.5. However, we find that depending on the dataset, base AutoFE algorithm used, and degree of non-IID across clients, the best performing sparsity on the feature mask greatly differs from 0.5. Thus, we have an outer loop that gradually refines the best performing sparsity.
Figure 8: Workflow of Algorithm 5
We provide a simple experiment on one of the datasets studying the efficacy of this federated feature selection technique. We use OpenFE on the jungle chess dataset. OpenFE is known to create an extremely large number of features, so this is a good candidate AutoFE method to test our federated feature selection method. We perform OpenFE on each client separately then find the union of the engineered features. We perform 3 trials in the IID case (each of these trials takes a very long time as the new feature space is so large with OpenFE that evaluations for hyperparameter tuning take a very long time). The mean (std dev) base test score with raw features is 0.6804 (0.0041). For the engineered features without federated feature selection it is 0.6349 (0.0860). For the engineered features with federated feature selection it is 0.7654 (0.0104). This represents -6.69% change without federated feature selection and a 12.49% change with federated feature selection. This shows that the federated feature selection we propose in Algorithm 5 works quite well for AutoFE algorithms that generate many features (in the case of OpenFE with jungle chess dataset across 8 clients, there were around 850 generated features in the union).